OpenStudy (babyslapmafro):

Could someone please point out my error solving the following problem involving a cross and dot product. Find u (dot) (v x w) u=<2,-3,1> v=<4,1,-3> w=<0,1,5> My answer=-40 Answer in book=80

OpenStudy (babyslapmafro):

v x w = <8,20,4> u (dot) <8,20,4>=<2,-3,1> (dot) <8,20,4>=-40

OpenStudy (anonymous):

$$\vec{v}\times\vec{w}=(5+3,0-20,4-0)=(8,-20,4)\\\vec{u}\cdot(\vec{v}\times\vec{w})=(2,-3,1)\cdot(8,-20,4)=16+60+4=80$$

OpenStudy (anonymous):

You messed up the sign of your middle component when computing your cross product!

OpenStudy (babyslapmafro):